DIRC Operations Report

Monday 2007/12/03 to Wednesday 2007/12/12


Oncall: Rolf Andreassen (Jose Benitez on backup)
Fast Monitoring: Rolf Andreassen, Jose Benitez

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: 0 x 1034 (last week: 0 x 1034)

PEP delivery: 0 fb-1 (last week: 0 fb-1)

BaBar integrated lumi:0 fb-1 (last week: 0 fb-1)

Run 7 total delivery: 0 fb-1

Run 7 total BaBar recorded: 0 fb-1 total, including 0 fb-1 off-peak data

PEP activities/issues since last DIRC meeting

  • Whole period: Warm up accelerator for restart.
  • Dec 12th: Electrons in the North Damping Ring.

    BaBar activities/issues since last DIRC meeting

  • Take cosmics and calibrations to make the machine limber.

    DIRC activities/issues since last DIRC meeting

  • Dec 7th to Dec 11th: Nitrogen flows fluctuated quite a bit in Epics, causing alarms. We first opened the alarm thresholds for sectors 0 and 10. On Dec 11th we had a controlled access of Jerry, Matt and Jose to look at the electronics; they found nothing obviously wrong, but the fluctuations went away while they were probing the boards.
  • Dec 10th: Some cryo work caused nitrogen flow to go to 0 for a few minutes. It came back quickly.
  • Dec 10th to 12th: Investigating the chiller glitches. They appear to be temperature-related; Matt sees a correlation with the glitches appearing whenever the hall temperature drops below the set point for the heater. This problem may be related to a relay, which we will replace. In the meantime the alarms have had to be masked.

    See DIRC logbook for details

    Lumi and Background since last DIRC meeting

    No background plot this week, since there are no beams.